# Comfortability of quantum walks on embedded graphs on surfaces

### v1

@@
Abstract. In this paper, a quantum walk model which reflects the underlying embedding on the surface is proposed. We obtain the scattering matrix of this quantum walk model characterized by the faces on the surface, and find a detection of the orientablility of the underlying embedding by the scattering information. The comfortability is the square norm of the stationary state restricted to the internal and reflected by the underlying embedding. We find that quantum walker feels more comfortable to a surface with small genus in some natural setting.
@@ 1 Introduction
How is the underlying geometric structure estimated by a reaction to an input? To tackle this inverse problem, we adopt a discrete-time quantum walk model, and try to extract a topological structure of graphs from the behavior of the quantum superposition. Among a lot of discrete-time quantum walk models [11], the Grover walk [19] may come to mind first. In fact, using the commutativity of the Grover matrix with any permutation matrix, the time evolution of the Grover walk is easily constructed [12], which exhibits interesting behaviors [9, 10] and also plays the key roles in the quantum search [17]. Moreover there are several important properties in the Grover walk connecting to not only a random walk[QCcoupling2, 2] but also the stationary Shrödinger equation [7].
Here, let us first apply the Grover walk as a discrete-time quantum walk model and focus on the embedding of graphs on surfaces as a geometric structure. For example, the complete graph with $4$ vertices $K_{4}$ has $11$ kinds of embedding way; see Figure 7. We expect that the quantum walk “feels” these underlying embeddings. Grover walk may decide whether an underlying graph admits a topological embedding on a given surface [3]. On the other hand, the time evolution of Grover walk depends only on the adjacent structure of the underlying graph because the time evolution operator $U_{Grover}$ acts as
@@
for any standard base labeled by an arc $e$ of the underlying graph. Here $o(e)$ and $t(e)$ are origin and terminal vertices of the arc $e$, and $\mathrm{deg}(x)$ is the degree of the vertex $x$. Then the weights associated with transmitting and reflecting at vertex $x$ are $2/\mathrm{deg}(x)$ and $12/\deg(x)-$ 1 in the Grover walk, which is independent of the underlying embedding. Therefore the Grover walk can not distinguish any two embedded surfaces of a graph. Thus for our object, we need to propose another kind of discrete-time quantum walk model for this problem.
To this end, let us introduce the idea of the drawing a graph without crossings of edges and extra faces on some closed surface, which is called the two-cell embedding [6, 14, 15]. More details are given in Section 3. See also Figures 5, 6 and 7. It is well known that the two-cell embedding on a closed surface of a graph $G=(X,E)$ is realized by the rotation system [6, 14, 15]. Here $X$ and $E$ are the set of vertices and the set ofunoriented edges. The rotation system is the triple of the symmetric digraph $G=(X,A)$, the rotation $\rho:A\to A$ and the twist $\tau:A\to\mathbb{Z}_{2}$, where $A$ is the set of the symmetric arcs induced by $E$. Here the rotation $\rho$ is decomposed into cyclic permutations with respect to the incoming arcs of each vertex. Thus our target can be switched to how a quantum walk model is constructed for given rotation system $(G,\rho,\tau)$.
In this paper, the abstract graph $G$ is deformed so that information of $\rho,\tau$ are reflected and also the in- and out- degrees are equally $2$. The regularity of the degree $2$ derives from the implementation by some optical polarizers elements [13]. See Sections 4.1, 4.2 for more detailed construction and also Figure 8. Such a new graph $G(\rho,\tau)$ is obtained by replacing each vertex of the double covering graph [6, 14] induced by $\tau$ with a directed cycle induced by $\rho$. The replaced directed cycles are called islands, and the symmetric arcs between the islands, which reflects the original adjacent relation, are called the bridges. The tails, which are semi-infinite paths, are inserted into all the island arcs. Such an assignment of tails is called the hedgehog; see Figure 8 (d). Let us explain the time evolution below; see Sections 4.3 and 4.4 for more detail. The degree-$2$ regularity of the new graph makes it possible that the one-step time evolution on the whole space can be described by the local scattering at each vertex of the internal graph by the $2\times 2$ unitary matrix
@@
Here $a,b,c,d$ are the complex valued weights associated with moving of a quantum walk for one step “from an island to the same island”, “ a bridge to an island”, “an island to a bridge” and “a bridge to the inverse bridge”, respectively. This unitarity of the local coin matrix $C$ includes also the unitarty of the total time evolution operator. On the other hand, we set that the dynamics on the tail is free; see (4.16) for the detailed dynamics of the free. The initial state is set so that the internal graph receives the constant inflow at every time step. If a quantum walker goes out to the external, it never come back to the internal because of the free dynamics of the tails. Such a quantum walker can be regarded as the outflow. By the balance between the in- and out- flows, this quantum walk model converges to a fixed point in the long time limit [4, 5, 8].
Now, for a fixed abstract graph $G_{o}$, let us vary the embedding that underlies it. For every $\rho$ and $\tau$, set $\psi_{\infty}$ as the stationary state and $\psi_{\infty}|_{G_{o}}$ as the restriction of $\psi_{\infty}$ to the internal graph $G_{o}$. In this paper, we focus on extracting embedding structures of a graph from the stationary state of this quantum walk model. To this end, we divide the stationary state into external and internal parts; we characterize the external part by the scattering matrix, which shows a reaction of the internal graph to an input at the tails, while the internal part by the comfortability, which is corresponding to the energy in the internal graph. We estimate how quantum walker
(i) gives a scattering on the underlying embedding by finding the expression of the scattering matrix
@@
that is, the larger $\mathcal{E}$ is, the more comfortable quantum walker feels to the underlying embedding.
Throughout this paper, in addition to the hedgehog tail condition, we subject the following assumption (2).
@@
    the assignment of tails is the hedgehog;
@@
Under such a construction of the quantum walk reflecting the underlying embedding on a closed surface, in this paper, now we are ready to state our main result on the scattering and comfortablity.
@@
Assume $d\in\mathbb{R}$ and set $\omega=-\det C$ and the hedgehog assignment of tails. Let $F$ be the set of faces induced by the rotation system $(G,\rho,\tau)$. The scattering matrix is decomposed into the following $|F|$ unitary matrices as follows:
@@
Here the operators induced by each face $f$, $I_{f}$ is the identity operator and $P_{f}$ are defined in (6.26).
Theorem 6.1 in Section 6 gives the scattering matrix in a more general setting. The operator $P_{f}$ is a weighted permutation matrix induced by the closed walk along the boudary of the face $f$, which is called the facial walk of $f$. The rotation system $(G,\rho,\tau)$ determines the set of faces $F$ and gives uniquely orientation of the closed walk along the boundary of the face $f\in F$. This closed walk corresponds to the facial walk, which is expressed by a sequence of arcs, and we simply write $f$ for this closed walk. We should remark $\bar{f}\not\in F$ for $f\in F$, where $\bar{f}$ is the inverse direction of $f$. The weight is determined by each edge type which is passed through the facial walk. Then using this property of the scattering matrix, the detection of the orientability by the scattering information is proposed in Theorem 6.2.
@@
Let $G=(X,A)$ and $(G,\rho,\tau)$ be the abstract symmetric digraph and its rotation system. Let $F$ be the set of faces determined by the rotation system $(G,\rho,\tau)$. Pick a tail at random from the hedgehog tails as an input. Under Assumption 1 with $a>0$ and $\omega=1$, the average of the comfortability with respect to a randomly chosen input is expressed by
@@
Here $f\cap\bar{f}$ is the set of self-intersection of the face $f\in F$; where if a facial walk $f$ passes through an arc $e$ and also its inverse $\bar{e}$, then the face $f$ is said to have a self-intersection at the unoriented boundary edge $|e|$.
Thus, the larger the first term, the more comfortable the quantum walker feels, while the larger the second term, the more uncomfortable it feels. The first term is related to an integer partition of $|A|$, while the second term is related to the self-intersections of faces. See Figure 4 for the self-intersection. Figure 2 shows the ranking of the comfortability of the embeddings for the complete graph $K_{4}$ with $a=0.98$. We discuss the messages of combinatorial structures from Theorem 1.2 in Section 2.
This paper is organized as follows. Section 2 is devoted to how geometric and combinatorial information of the graph embeddings is extracted from Theorem 1.2 as its corollaries. We show the best and worst embedding of the complete graph $K_{n}$ for quantum walker and also characterize the comfortability on the island by the Young diagram. In section 3, we give a short review on the graph embeddings on the orientable/non-orientable surfaces. In section 4, the time evolution of the quantum walk induced by the rotation system $(G,\rho,\tau)$ is discussed. In section 5, we show the unitary equivalence of the time evolutions between the isomorphic embeddings. In section 6, we find that the scattering matrix is characterized by the resulting faces and described by direct sum of unitary and circurant matrices. Moreover a detection method of the orientability by using the scattering information is proved. In section 7, we give the proof of Theorem 1.2.
Figure 1: The list of embeddings of $K_{4}$: The genus is described by $g$ and $k$, for orientable and non-orientable surfaces, respectively. The boundary lenghts of faces of the resulting embedding are indicated by $[\lambda_{1},\lambda_{2},\dots,\lambda_{\kappa}]$. For example, $[6,3,3]$ indicates there are $1$ hexagon and $2$ triangles in the embedding.
Figure 2: The ranking of the comfortability for the embeddings of $K_{4}$ with $a=0.98$: This ranking follows Corollary 2.1 in the setting of $a\to 1$.
@@
In this section, we discuss how more detailed geometric information is extracted from Theorem 1.2.
@@ 2.1 Observation 1: In the limit of $a\to 1$
On the other hand, if $a\to 1$, then the comfortability diverges. Then, taking $a=1-\delta$, we will find the appropriate scaling with respect to $\delta$, and its coefficient of the first order hoping that we will find some geometric information in the coefficient. Indeed, we obtain the following.
@@
The small genus increases the number of faces, while large genus induces longer faces and, therefore, larger number of intersections. Thus for $a\to 1$, quantum walker feels more comfortable to the surface with a smaller genus. In particular, the single-face embedding on an orientable surface attains the above comfortablity 00. Figure 2 shows the ranking of the comfortability for $K_{4}$ following Corollary 2.1.
@@
the first $\text{the first term}=\frac{1}{\delta^{2}}\left(\frac{|F|}{|E|}+\frac{|F|}{|E|}\delta+O(\delta^{2})\right),$
@@
$\frac{a}{|b|^{2}}=\frac{1}{2\delta}\left(1-\frac{1}{2}\delta+(\delta^{2})\right)$
@@
Then we have the desired conclusion. ∎
@@
For a fixed abstract graph, if the underlying embedding gives the most comfortability in all the possible embeddings of the graph, then it is called the best embedding of the graph, while if it gives the least comfortability, then it is called the worst embedding of the graph. Let us find the best and worst embeddings of the complete graph $K_{n}$ by using the following famous graph theoretical facts.
@@
    The (minimal) genus of $K_{n}$  (Rindel and Youngs (1968) [18]):
@@
By combining Corollary 2.1 with Fact 1, the most comfortable underlying embedding of $K_{n}$ for quantum walker can be chractorized as follows.
@@
Under the setting of Corollary 2.1, the best and worst embeddings of the complete graph $K_{n}$ for the comfortability of quantum walker must be on the closed surfaces with the minimal and maximal genus, respectively, whose orientability are divided into cases of $\mod(n,4)$ as follows.
@@
    $n\equiv 1,2\mod 4$ case:
    • The best embedding: any non-orientable surface
    • The worst embedding: any orientable surface
@@
    $n\equiv 0,3\mod 4$ and $n\neq 3,4,7$ case:
    • The best embedding: any orientable and non-orientable surfaces
    • The worst embedding: a non-orientable surface
@@
    $n=3,4,7$ case:
    • The best embedding: any orientable surface
    • The worst embedding: a non-orientable surface
The best and worst surfaces of $K_{n}$ for quantum walker is described in the following Table.
$n\equiv 1,2\mod 4$ $n\equiv 0,3\mod 4$, $n\neq 3,4,7$ $n=3,4,7$
Best Non-ori Ori and Non-ori Ori
Worst Ori Non-ori Non-ori
Table 1: The best and worst surfaces of $K_{n}$ for quantum walker: The genera for the best and worst surfaces are minimal and maximal genera, respectively. The best embedding is close to the triangulation, because $2|E|=3(|E|-|V|+2-(n-3)(n-4)/6)$ holds. In particular, if $n\equiv 0,3\mod 4$ with $n\not\equiv 8,11\mod 12$, then the best embedding is the triangulation. Since every worst embedding is a single-face embedding, the comfortability of the worst embedding is 00 for $n\equiv 1,2\mod 4$.
Proof.
Let us find that the best embedding and the worst embedding of $K_{n}$ for quantum walker as follows.
  1. (1)
    Best: The expression (2.2) tells us that the large number of faces (the first term) and small number of self-intersection (the second term) makes quantum walker feel comfortable. Thus we expect that the small genus embedding will be the best embedding because the small genus accomplishes both of them.
    First let us estimate the number of faces for the minimum genus embedding. Combining Fact 1 with the Eular formula
    $|F|=|E|-|V|+2-\begin{cases}2\gamma(K_{n})&\text{: Orientable case}\\ \tilde{\gamma}(K_{n})&\text{: Non-orientable case}\end{cases},$
@@
    there are no intersections in the minimum genus embeddings on both “non-orientanble surfaces” and “orienable surfaces except $n\equiv 2,5\mod 12$”.
    Put $n$, $m=n(n-1)/2$ and $\ell$ as the numbers of vertices, edges and faces for the resulting embbeding of $K_{n}$, respectively. Assume that there is a face having a self-intersection in an embedding of $K_{n}$. Note that the boundary length of a face having a self-intersection must be at least $8$. Then the following inequality holds.
@@
      Non-orientable case: Let $k$ be the genus of the underlying closed surface of the embedding having self-intersections. By the Euler formula and (2.5), we have
@@
      and the embeddings on the non-orientable surfaces with the minimal genus has no self-intersections. Then the second term of (2.2) for the minimal genus embedding on the non-orientable surface is reduced to 00.
@@
      except $n\equiv 2,5\mod 6$. Thus it is ensured that there are no interactions in the minimum genus embedding except $n\equiv 2,5\mod 12$. So the second term of (2.2) for the minimal genus embedding on the orientable surface is reduced to 00 except $n\equiv 2,5\mod 12$.
@@
    Worst: The expression (2.2) tells us that the small number of faces (the first term) and large number of self-intersection (the second term) makes quantum walker feel uncomfortable. Thus we expect that the large genus embedding will be the worst embedding because the large genus accomplishes both of them.
@@
      $n\equiv 1,2\mod 4$ case: If the underlying surface is orientable, then there is only one face and every boundary face has the self-intersection, which means that comfortability is reduced to 00 by (2.2). On the other hand, if the underlying surface is non-orientable, then there must exists at least one edge having no self-intersection in the face, which implies that the comfortability is non-zero by (2.2). Thus we have the worst embedding is the maximal genus embedding on the orientable surface.
@@
      $n\equiv 0,3\mod 4$ case: If the underlying surface is orientable, then the number of faces is $2$. Then the boundary edges of the two faces have no self-intersections. If a twist is inserted into one of the edges of the boundary edges so called the edge twisted surgery [6, 14], then the two faces are merged into a single face conserving the self intersected edges and the resulting surface becomes non-orientable. Then this single-face embedding on the non-orientable surface is worse than the embedding with double-face on the orientable surface.
@@
Therefore (2.10) and (2.11) lead the desired conclusion. ∎
@@
It is shown in the proof of Theorem 7.1 that the comfortability on the island $\mathcal{E}|_{A_{is}}$ is proportional to the first term of (1.1). Then let us estimate the first term of (1.1) by using some combinatorial method.

### v2

@@
Abstract. The time evolutions of discrete-time quantum walks on graphs are determined by the local adjacency relations of the graphs. In this paper, first, we construct a discrete-time quantum walk model that reflects the embedding on the surface so that an underlying global geometric information is reflected. Second, we consider the scattering problem of this quantum walk model. We obtain the scattering matrix characterized by the faces on the surface and detect the orientablility of the embedding using scattering information. For the stationary state in the scattering problem, the comfortability is defined as the square norm of the stationary state restricted to the internal. This indicates how a quantum walker is stored in the internal under the embedding. Then we find that a quantum walker feels more comfortable on a surface with small genus in some natural setting. We illustrate our results with some interesting examples.
@@ 1 Introduction
For the scattering problems of finite graphs embedded on the surfaces, how is the underlying geometric structure estimated by a reaction to an input? To tackle this problem, we adopt a discrete-time quantum walk model and attempt to extract topological structures of graphs from the behavior of quantum superposition. Among a lot of discrete-time quantum walk models [15], the Grover walk [24] may come to mind first. In fact, using the commutativity of the Grover matrix with any permutation matrix, the time evolution of the Grover walk is easily constructed [16], which exhibits interesting behaviors [11, 14] and also plays a key role in the quantum search [21]. Moreover, there are several important properties of the Grover walk that connect not only to a random walk[3, 2] but also to the stationary Shrödinger equation [8].
Here, let us first apply the Grover walk as a discrete-time quantum walk model and focus on the embedding of graphs into closed surfaces $\mathbb{F}^{2}$ as a geometric structure. Note that any finite graph can be embedded in $3$-dimensional Euclidean space $\mathbb{R}^{3}$. For example, the complete graph with $4$ vertices $K_{4}$ has $11$ embeddings up to the homeomorphisms of the surfaces; see Figure 7. We expect that the quantum walk “feels” these embeddings. Grover walk may decide whether an underlying graph admits a topological embedding on a given surface in a similar manner to the results in [4]. On the other hand, the time evolution of Grover walk depends only on the adjacency structure of the underlying graph because the time evolution operator $U_{Grover}$ acts as
@@
for any standard base labeled by arc $e$ of the underlying graph. This shows the local scattering at the vertex $x$ of the incident wave represented by the arc $e$ [9, 23]. Here $o(e)$ and $t(e)$ are the origin and terminal vertices of the arc $e$, $\bar{e}$ is the inverse arc of $e$ and $\mathrm{deg}(x)$ is the degree of the vertex $x$. Then, the weights associated with transmitting and reflecting at the vertex $x$ are $2/\mathrm{deg}(x)$ and $12/\deg(x)-$ 1 in the Grover walk, which is independent of the embedding. Therefore, the Grover walk cannot distinguish between any two embedded surfaces of a graph. Thus, for our objective, we propose another type of discrete-time quantum walk model.
To this end, we use the concept of drawing a graph without crossings of edges and extra faces on a closed surface, which is called the two-cell embedding [7, 18, 19]. More details are given in Section 3. See also Figures 5, 6 and 7. It is well known that two-cell embedding on the closed surface of a graph $G=(X,E)$ can be realized using a rotation system [7, 18, 19], where $X$ and $E$ are sets of vertices and unoriented edges, respectively. We put $A$ as the set of symmetric arcs induced by $E$. The rotation system is the triple of the symmetric digraph $G=(X,A)$, the rotation $\rho:A\to A$ and the twist $\tau:A\to\mathbb{Z}_{2}$. Here the rotation $\rho$ is decomposed into cyclic permutations with respect to the incoming arcs of each vertex. Thus our target can be switched to constructing a quantum walk model for a given rotation system $(G,\rho,\tau)$. We have proposed a quantum walk model on the orientable surfaces in [12] and characterized the stationary state by some spanning subgraphs of the dual graph induced by the orientable embeddings. Here we will propose an extension model of [12] that can be constructed on both orientable and non-orientable surfaces as follows. In this paper, the abstract graph $G$ is deformed so that the information of $\rho,\tau$ is reflected, and the in- and out-degrees are equal $2$. The regularity of degree $2$ is derived from the implementation of optical polarizer elements [17]. See Sections 4.1 and 4.2 for a more detailed construction and Figure 8. Such a new graph $G(\rho,\tau)$ is obtained by replacing each vertex of the double covering graph [7, 18] induced by $\tau$ with a directed cycle induced by $\rho$. The replaced directed cycles are called islands, and the symmetric arcs between the islands that reflect the original adjacent relation, are called the bridges. The tails, which are semi-infinite paths, are inserted into all the island arcs. Such an assignment of tails is called the hedgehog; see Figure 8 (d). Let us explain the time evolution below; see Sections 4.3 and 4.4 for further details. The degree-$2$ regularity of the new graph makes it possible that the one-step time evolution on the whole space can be described by the local scattering at each vertex of the internal graph by the $2\times 2$ unitary matrix
@@
Here $a,b,c,d$ are the complex valued weights associated with moving of a quantum walk for one step “from an island to the same island,” “ a bridge to an island,” “an island to a bridge” and “a bridge to the inverse bridge,” respectively. The unitarity of the local coin matrix $C$ also includes the unitarity of the total time evolution operator. On the other hand, we set that the dynamics on the tail is free; see (4.19) for the detailed dynamics of the free. The initial state is set so that the internal graph receives constant inflow at every time step. If a quantum walker goes outside the interior, it never returns to the interior, owing to the free dynamics of the tails. Such a quantum walker can be regarded as outflow. By balancing the in- and out-flows, this quantum walk model converges to a fixed point in the long time limit [5, 6, 10].
Now, for a fixed abstract graph $G_{o}$, let us vary its embedding onto the surfaces. For every $\rho$ and $\tau$, set $\psi_{\infty}$ as the stationary state and $\psi_{\infty}|_{G_{o}}$ as the restriction of $\psi_{\infty}$ on the internal graph $G_{o}$. In this study, we focus on extracting the embedding structures of a graph from the stationary state of this quantum walk model. To this end, we divide the stationary state into external and internal parts; we characterize the external part by the scattering matrix, which shows the reaction of the internal graph to an input at the tails, while the internal part by the comfortability, which corresponds to the energy in the internal graph. We estimate how quantum walker
(i) gives a scattering on the embedding by finding the expression of the scattering matrix
@@
that is, the larger $\mathcal{E}$ is, the more comfortable quantum walker feels to the embedding.
Throughout this paper, in addition to the hedgehog tail attachment, we impose the following assumption (2).
@@
    the assignment of tails is the hedgehog, more precisely see (6.34);
@@
Under such a construction of the quantum walk reflecting the embedding on a closed surface, now we are ready to state our main result on the scattering and comfortablity.
@@
Assume $d\in\mathbb{R}$ and set $\omega=-\det C$ and the hedgehog assignment of the tails. Let $F$ be the set of faces induced by the rotation system $(G,\rho,\tau)$. The scattering matrix is decomposed into the following $|F|$ unitary matrices as follows:
@@
Here $I_{f}$ is the identity matrix and $P_{f}$ is a unitary and weighted permutation matrix corresponding to the closed walk on the boundary of face $f$ defined in (6.35).
Theorem 6.1 in Section 6 gives the scattering matrix in a more general setting. The operator $P_{f}$ is a weighted permutation matrix induced by the closed walk along the boundary of face $f$, which is called the facial walk of $f$. The rotation system $(G,\rho,\tau)$ determines the set of faces $F$ and provides a unique orientation for the closed walk along the boundary of face $f\in F$. This closed walk corresponds to the facial walk, which is expressed by a sequence of arcs, and we simply write $f$ for this closed walk. It should be noted that $\bar{f}\not\in F$ for $f\in F$, where $\bar{f}$ is the inverse direction of $f$. The weight is determined by each edge type that passes through the facial walk. Based on this property of the scattering matrix, detection of orientability using scattering information is proposed in Theorem 6.2.
It is shown in Proposition 6.1 that the scattering matrices of two equivalent embeddings are unitarily equivalent. Although the scattering of the two embeddings with the same inflow is generally different, the comfortability $\mathbb{E}(\mathcal{E})$ under the following setting is invariant under the isomorphism of the embeddings. The reason for this is explained in Section 7.
@@
Let $G=(X,A)$ and $(G,\rho,\tau)$ be the abstract symmetric digraph and its rotation system, respectively. Let $F$ be the set of faces determined by the rotation system $(G,\rho,\tau)$. Randomly pick a tail from the hedgehog tails as an input. Let $\mathbb{P}_{1},\dots,\mathbb{P}_{m}$ be the hedgehog tails. Suppose $\boldsymbol{\alpha}_{in}$ is a random variable distributed uniformly in $\{\delta_{j}\;|\;j=1,\dots,m\}$. Under Assumption 1 with $a>0$ and $\omega=1$, the average of the comfortability with respect to a randomly chosen input is expressed by
@@
Here $f\cap\bar{f}$ is the set of self-intersections of the face $f\in F$, where if a facial walk $f$ passes through an arc $e$ and also its inverse $\bar{e}$, then the face $f$ is said to have a self-intersection at the unoriented boundary edge $|e|$.
Thus, the larger the first term is, the more comfortable the quantum walker feels, whereas the larger the second term is, the more uncomfortable the walker feels. The first term is related to an integer partition of $|A|$, whereas the second term is related to self-intersections of faces. See Figure 4 for the self-intersection. Figure 2 shows the ranking of the comfortability of the embeddings for the complete graph $K_{4}$ with $a=0.98$. We discuss the messages of combinatorial structures from Theorem 1.2 in Section 2.
The remainder of this paper is organized as follows. Section 2 describes how the geometric and combinatorial information of graph embeddings can be extracted from Theorem 1.2 as corollaries. We show the best and worst embeddings of the complete graph $K_{n}$ for a quantum walker and also characterize the comfortability on the island using the Young diagram. In Section 3, we provide a short review of graph embeddings on orientable/non-orientable surfaces. In Section 4, we describe the time evolution of the quantum walk induced by the rotation system $(G,\rho,\tau)$. In Section 5, we present the unitary equivalence of the time evolutions between the equivalent embeddings. In Section 6, we show that the scattering matrix is characterized by the resulting faces and is described by the direct sum of the unitary and circulant matrix matrices. Moreover, a method for detecting orientability using scattering information is proved. In Section 7, we discuss the comfortability and give the proof of Theorem 1.2.
Figure 1: The list of embeddings of $K_{4}$: The genus is described by $g$ and $k$, for orientable and non-orientable surfaces, respectively. For example, $g=0$ and $g=1$ correspond to the surfaces of the sphere and torus, respectively, whereas $k=1$ and $k=2$ correspond to the surfaces of the projective plane and Klein’s bottle, respectively. The boundary lengths of the faces of the resulting embedding are denoted as $[\lambda_{1},\lambda_{2},\dots,\lambda_{\kappa}]$. For example, $[6,3,3]$ indicates that one hexagon and two triangles exist in the embedding.
Figure 2: The ranking of the comfortability for the embeddings of $K_{4}$ with $a=0.98$: This ranking follows Corollary 2.1 in the setting of $a\to 1$. The most comfortable embedding for $K_{4}$ is the embedding on the sphere; the worst is the embedding on the non-orientable surface with the maximum genus $k=3$.
@@
In this section, we discuss the extraction of more detailed geometric information from Theorem 1.2.
@@ 2.1 Observation 1: In the limit of $a\to 1$
On the other hand, if $a\to 1$, then the comfortability diverges. Then, taking $a=1-\delta$, we will find the appropriate scaling with respect to $\delta$, and its coefficient of the first order. We hope that the coefficient can be characterized using geometric information. Indeed, we obtain the following.
@@
The smaller genus increases the number of faces, whereas the larger genus induces larger faces and then increases the number of self-intersections. Thus, for $a\to 1$, the quantum walker feels more comfortable on the surface with a smaller genus. In particular, the single-face embedding on an orientable surface attains the above comfortablity 00. Figure 2 shows the ranking of the comfortability for $K_{4}$ following Corollary 2.1.
@@
the first $\text{the first term}=\frac{1}{\delta^{2}}\left(\frac{|F|}{|E|}+\frac{|F|}{|E|}\delta+O(\delta^{2})\right).$
@@
$\frac{a}{|b|^{2}}=\frac{1}{2\delta}\left(1-\frac{1}{2}\delta+(\delta^{2})\right),$
@@
Thus, we arrive at the desired conclusion. ∎
@@
For a fixed abstract graph, if the embedding gives the most comfortability in all the possible embeddings of the graph, then it is called the best embedding of the graph, whereas if it gives the least comfortability, then it is called the worst embedding of the graph. Let us determine the best and worst embeddings of the complete graph $K_{n}$ using the following famous graph theoretical facts:
@@
    The (minimal) genus of $K_{n}$  (Ringel and Youngs (1968) [22]):
@@
By combining Corollary 2.1 with Fact 1, the most comfortable embedding of $K_{n}$ for quantum walker can be characterized as follows.
@@
Under the setting of Corollary 2.1, the best and worst embeddings of the complete graph $K_{n}$ are as follows:
@@
    If $n\equiv 1,2\mod 4$,
    • the best embedding is
      any embedding on the non-orientable surface with minimal genus;
    • the worst embedding is
      any embedding on the orientable surface with maximal genus.
@@
    If $n\equiv 0,3\mod 4$,
    • the best embedding is
      any embedding on the orientable and the non-orientable surfaces with minimal genus for $n\neq 3,4,7$;
      any embedding on an orientable surface with a minimal genus for $n=3,4,7$;
    • the worst embedding is
      embedding on a non-orientable surface with a maximum genus.
The best and worst surfaces of $K_{n}$ for the quantum walker are listed in the following table.
$n\equiv 1,2\mod 4$ $n\equiv 0,3\mod 4$, $n\neq 3,4,7$ $n=3,4,7$
Best (the minimal genus) Non-ori Ori and Non-ori Ori
Worst (the maximal genus) Ori Non-ori Non-ori
Table 1: The best and worst surfaces of $K_{n}$ for quantum walker: The genera for the best and worst surfaces are minimal and maximal genera, respectively. The best embedding is close to the triangulation, because $2|E|=3\{|E|-|V|+(2-(n-3)(n-4)/6)\}$ holds, where the right hand side is “almost” $3|F|$, see (2.4). In particular, if $n\equiv 0,3\mod 4$ and $n\not\equiv 8,11\mod 12$, then every triangulation is the best embedding. Since every worst embedding is a single-face embedding, the comfortability of the worst embedding is 00 for $n\equiv 1,2\mod 4$.
Proof.
Let us find that the best embedding and worst embedding of $K_{n}$ for a quantum walker are as follows.
@@
    Best: The expression (2.2) indicates that a large number of faces (the first term) and a small number of self-intersections (the second term) make the quantum walker feel comfortable. Thus, we expect that a small genus embedding will be the best embedding because the small genus accomplishes both of them.
    First let us estimate the number of faces for minimum genus embedding. Combining Fact 1 with the Eular formula
    $|F|=|E|-|V|+2-\begin{cases}2\gamma(K_{n})&\text{: Orientable case,}\\ \tilde{\gamma}(K_{n})&\text{: Non-orientable case,}\end{cases}$ (2.4)
@@
    there are no self-intersections in the minimum genus embeddings on both “non-orientanble surfaces” and “orienable surfaces except $n\equiv 2,5\mod 12$”.
    Let $n$, $m=n(n-1)/2$ and $\ell$ be the numbers of vertices, edges, and faces for the resulting embedding of $K_{n}$, respectively. Assume that there is a face with self-intersection in the embedding of $K_{n}$. Note that the boundary length of a face with a self-intersection must be at least $8$. Then the following inequality holds.
@@
      Non-orientable case: Let $k$ be the genus of the underlying closed surface of an embedding with self-intersections. By the Euler formula and (2.6), we have
@@
      and embeddings on non-orientable surfaces with the minimal genus have no self-intersections. Subsequently, the second term of (2.2) for the minimal genus embedding on the non-orientable surface is reduced to 00.
@@
      with the exception of $n\equiv 2,5\mod 12$. Thus, it is ensured that there are no interactions in the minimum genus embedding, except $n\equiv 2,5\mod 12$. Thus, the second term of (2.2) for minimal genus embedding on the orientable surface is reduced to 00 except $n\equiv 2,5\mod 12$.
@@
    Worst: The expression (2.2) tells us that the small number of faces (the first term) and the large number of self-intersections (the second term) make the quantum walker feel uncomfortable. Thus we expect that the large genus embedding will be the worst embedding because the large genus accomplishes both of them.
@@
      $n\equiv 1,2\mod 4$ case: If the underlying surface is orientable, then there is only one face, and every boundary face has a self-intersection, which means that comfortability is reduced to 00 by (2.2). On the other hand, if the underlying surface is non-orientable, then there must exist at least one edge with no self-intersection in the face, which implies that the comfortability is non-zero by (2.2). Thus the worst embedding is the maximal genus embedding on the orientable surface.
@@
      $n\equiv 0,3\mod 4$ case: If the underlying surface is orientable, then the number of faces is $2$. Then the boundary edges of the two faces have no self-intersections. If a twist is inserted into one of the edges of the boundary edges so called the edge twisted surgery [7, 18], then the two faces are merged into a single face, conserving the self-intersected edges, and the resulting surface becomes non-orientable. This single-face embedding on the non-orientable surface is worse than the embedding with double-face on the orientable surface.
@@
Therefore, (2.11) and (2.12) lead to the desired conclusion. ∎
@@
The comfortability on the island $\mathcal{E}|_{A_{is}}$ is proportional to the first term of (1.1), which will be shown in the proof of Theorem 7.1. Then, let us estimate the first term of (1.1) using a combinatorial method.